5 times a number is 14 less than the square of that number. Find the negative solution.

Answer by ikleyn(52781)

.

5x = x^2 - 14


x^2 - 5x - 14 = 0


(x-7)*(x+2) = 0


Of the two roots,  x= 7  and  x= -2,  only one is negative  x= -2.


ANSWER.  -2.

Solved.